Operation Summary for February 15, 2010 - February 17, 2010 |
Monday Feb. 15 The day shift began with Operations monitoring the Tevatron (TeV) store 7611 (144.07E30) and the Recycler stash (118.2E10), with the Antiproton Source stacking (12.4E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and Meson MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, Operations reported that Jasmin experimenters accessed M01 for foil removal. At 10:30 AM, Operations put the Proton Source into a one-hour downtime to allow Booster personnel to work on TCLK and for Preaccelerator personnel to flush the H- Source water resistor and replace its extractor. At 11:43 AM, a TeV RF station (TRF1) tripped off with no indication; 7E30 of luminosity was lost. Operators secured the MTA - MuCool enclosure. At 2:06 PM, the MTA beam testing started and lasted until the end of shift; 58 of the 60 allowed pulses were taken. MI personnel conducted coalescing studies. At the end of the shift the NuMI beam permit tripped on a fan differential pressure. NuMI personnel investigated. During the evening shift, MTA beam studies wrapped up early on the shift. NuMI personnel had their fan problems resolved by 8:16 PM. Operations resumed sending beam to NuMI. Operations disabled the MTest beam at 6 PM. Operations terminated store 7611 at 10:08 PM and put the TeV into shot setup. Operations established store 7612 at 11:47 PM with an Average Initial Luminosity (AIL) of 285.16E30. Tuesday Feb. 16 The midnight shift began with Operations monitoring store 7612 (251.57E30) and the stash (68.5E10), with Pbar stacking (11.6E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations tuned MI to improve beam positions. Operations reported a Linac RF station (LRF4) power amplifier crowbar, which caused some downtime. Operations established beam to MTest shortly after 4 AM. During the day shift, Operations prepared NuMI for an access at 9:25 AM. Linac personnel began the power amplifier replacement for LRF4 at 9:50 AM. The Tevatron Electron Lens tripped off on a vacuum interlock indication; it reset. Linac resumed sending beam at 12:30 PM. Jasmin experimenters accessed M01 to change out a foil. During the evening shift, Operations terminated store 7612 at 5:45 PM and put the TeV into shot setup. Operations disabled the MTest beam at 5:52 PM. Operations established store 7613 at 7:24 PM with an AIL of 297.24E30. Wednesday Feb. 17 The midnight shift began with Operations monitoring store 7613 (148.24E30) and the stash (156.3E10), with Pbar stacking (26.7E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, at 5:25 AM, a TeV RF station (TRF1) tripped off and caused a luminosity loss of 8E30. At 6:20 AM, the Feynman Computer Center lost power. The Plans for Wednesday The plans for today are to check out TRF1 in between stores and then continue with stacking, stashing, and storing. |
|
More Information Comments and Suggestions |
| last modified 02/19/2010 email Fermilab |
| Security, Privacy, Legal | |